A ball is thrown vertically upwards at a speed of 11.0 m s−1.
What is the maximum height it reaches?
1
Expert's answer
2020-11-19T09:22:45-0500

Solution

Using equation of motion

"v^2=u^2+2as"

"(0)^2=(11)^2-2gH"

At maximum height velocity is zero.

"H=\frac{121}{2\times9.8}=6.17m"